OpenStudy (anonymous):

after solving a system of linear equations, how do I check my answer to make sure that it is correct?

myininaya (myininaya):

pretend we have x+y=3 and y=2x-1 so x+2x-1=3 3x-1=3 3x=4 x=4/3 y=2x-1=2(4/3)-1=8/3-3/3=5/3 (x,y)=(4/3,5/3) to check plug in x to both and see if you get the same y for both equations so x=4/3 4/3+y=3 so y=3-4/3=9/3-4/3=5/3 y=2(4/3)-1=8/3-3/3=5/3 since we got the same y for both equations we are good

OpenStudy (anonymous):

ok thanks for the tip. I was completely confused before bu I am better now so thanks

myininaya (myininaya):

np